# Language Constraints on Language Models’ Inductive Inferences

### v1

@@ 1 Introduction
A hallmark of human cognition is our ability to make inductive inferences about categories (Osherson et al., 1990; Murphy, 2004). We can readily extend immediately available information along different categorical structures—e.g., on learning that robins have a novel property, we might extend it to other birds. An active area of research in cognitive science is to understand exactly how these inferences are constrained (Gelman et al., 2002; Hollander et al., 2002). For instance, a goldfish is a pet, a fish, an animal, and a living thing, and so deciding the extent to which properties generalize is a non-trivial task—one which we master anyway.
@@
Figure 1: We study how the surface form of a proposition constrains a model’s inductive inferences. We compare between universal quantifiers (all), generics (bare plurals), and indefinite quantifiers (some). Humans represent them differently and show consistent graded effects in their inductive inferences (Gelman et al., 2002).
@@
Overall, our experiments and results contribute to the broader literature surrounding (V)LMs and their conceptual knowledge. Existing work on models’ inductive reasoning capacities has largely focused on the effect of the conceptual content of the premises (Misra et al., 2021, 2022; Han et al., 2024; Bhatia, 2023). For instance, Han et al. (2024) use stimuli from Osherson et al. (1990) which do not consider the effect of scope-modifying cues on their stimuli. We contribute to this body of work by narrowing our focus on the constraints imposed by various surface form changes, keeping the conceptual content the same (i.e., same concepts). Furthermore, our work also complements recent work on generics and (V)LMs. For instance, recent work has focused on the extent to which these models are able to reason about generics in the context of their interplay with exceptions (Allaway et al., 2023, 2024; Frank and Allaway, 2025). While generics have been explicitly compared against quantifiers (Ralethe and Buys, 2022; Collacciani et al., 2024; Cilleruelo et al., 2025), no work has investigated the differences between generics and quantifiers in terms of their inductive constraints, which is an important component of their meaning. Finally, an unintended contribution of our work is our developmentally inspired benchmark focusing on models’ behavior on all and some, which we release in both the language and the vision+language modalities, and is the first of its kind, to the best of our knowledge—with previous work instead focusing on vague quantifiers (Wong et al., 2025). This joints other developmentally inspired benchmarks for VLMs (Tan et al., 2024; Yiu et al., 2025). We make our code available at https://github.com/kanishkamisra/inductive-constraints.
@@ 3.2 Results and Analysis
To recap our stimuli, each category is associated with 5 total images, and for each image, we have 4 types of questions—one positive question which asks for the presence of the category (expecting ‘Yes’), and three negatives, which ask for the presence of other category (expecting ‘No’). A competent model must be able to identify not only that the object in all images (per category) is the target category but also that it is not some other category. This results in two types of measures. First, for each question type, we measure the average percentage of cases the model correctly answers the question correctly for all images for a given category. We report this accuracy measure per question-type. Additionally, we also measure the average proportion of time the model correctly answers all four questions for all five images in the given category—i.e., all 20 questions associated with a given category. We refer to this latter measure ‘Joint’ accuracy. Table˜1 shows the accuracy measure across all question types as well as the Joint accuracy.
@@ 4.1 Data
Due to a lack of existing datasets that test knowledge of quantifiers in VLMs in a controlled manner, we design our own Vision and Language dataset, inspired by Gelman et al. (2002) and Smith (1980). To this end, we prompt Gemini 2.5 Flash Image (Comanici et al., 2025) to first generate an image with a specified property that holds true for either all or some objects (e.g., flowers in a vase), we then prompt it to edit the image by adding new objects such that the property now applied to the opposite quantifier (e.g., adding a few flowers outside the vase). This gives us our input context variation between all and some. To ensure the generated images are consistent with our instructions, we manually analyzed every generation. Specifically, we prompted the model to generate 200 pairs of images, and then checked if 1) the pairs were minimally different (no visible difference in backgrounds), still maintained the all vs. some distinction, and 3) did not add additional properties beyond what was specified. This resulted in 140 pairs in total. Appendix C shows full details about our manual annotation. We then paired these images with their respective questions, as described in the previous paragraph (6 different conditions), giving us a total of 840 stimuli. These stimuli mimic the kinds of stimuli used by Gelman et al. (2002).
@@
Our language stimuli include synthetically generated contexts posed as word-problems about objects and their properties. We generate 150 unique contexts differing in objects (N=5, e.g., coins in a jar, bottles in a create, etc.), property pairs (N=6, e.g., red/blue in color, contains milk/soda, etc.), and amounts (N=5, e.g., 500 total objects, with 300 having a given property and 200 having the other property). This results in a total of 900 stimuli (when combined with 6 different conditions). The example in Table 2 uses as its object ‘blocks on a table’, as its property pair, ‘red/blue in color’ and as its amount, 20 total objects—12 having the majority property (red) and 8 having the minority property (blue).
@@ C.1 Prompts
In the third condition, we first prompt the model to generate a image with most of the objects (majority) in a container, and some of them (minority) out of it. Then we prompt the model to modify the image by removing the objects that are out of the container.
@@ C.2 Annotation
We manually annotate the 200 generated image pairs using the frontend shown in Figure 6. We excluded 60 pairs from our Examples of excluded pairs are shown in Figure 7.

### v2

@@ 1 Introduction
A hallmark of human cognition is our ability to make inductive inferences about categories (Osherson et al., 1990; Murphy, 2004). We can readily extend immediately available information along different categorical structures—e.g., on learning that robins have a novel property, we might extend it to other birds. An active area of research in cognitive science is to understand exactly how these inferences are constrained (Gelman et al., 2002; Hollander et al., 2002; Cimpian and Markman, 2008). For instance, a goldfish is a pet, a fish, an animal, and a living thing, and so deciding the extent to which properties generalize beyond available evidence, to other categories, is a non-trivial task—one which we master anyway.
@@
Figure 1: We study how the form of a proposition constrains a model’s inductive inferences. We compare between universal quantifiers (all), generics (bare plurals), and indefinite quantifiers (some). Humans represent them differently and show consistent graded effects in their inductive inferences (Gelman et al., 2002).
@@
Overall, our experiments and results contribute to the broader literature surrounding (V)LMs and their conceptual knowledge. Existing work on models’ inductive reasoning capacities has largely focused on the effect of the conceptual content of the premises (Misra et al., 2021, 2022; Han et al., 2024; Bhatia, 2023). For instance, Han et al. (2024) use stimuli from Osherson et al. (1990), which does not contain any variation in terms of scope-modifying cues. We contribute to this body of work by narrowing our focus on the constraints imposed by various surface form changes, keeping the conceptual content the same (i.e., same concepts). Furthermore, our work also complements recent work on generics and (V)LMs. For instance, recent work has focused on the extent to which these models are able to reason about generics in the context of their interplay with exceptions (Allaway et al., 2023, 2024; Frank and Allaway, 2025). While generics have been explicitly compared against quantifiers (Ralethe and Buys, 2022; Collacciani et al., 2024; Cilleruelo et al., 2025), no work has investigated the differences between generics and quantifiers in terms of their inductive constraints, which is an important component of their meaning. Finally, an unintended contribution of our work is our developmentally inspired benchmark focusing on models’ behavior on all and some, which we release in both the language and the vision+language modalities, and is the first of its kind, to the best of our knowledge—with previous work instead focusing on vague quantifiers (Wong et al., 2025). This joins other developmentally inspired benchmarks for VLMs (Tan et al., 2024; Yiu et al., 2025). Our code is available at https://github.com/kanishkamisra/inductive-constraints.
@@ 3.2 Results and Analysis
To recap our stimuli, each category is associated with 5 total images, and for each image, we have 4 types of questions—one positive question which asks for the presence of the category (expecting ‘Yes’), and three negatives, which ask for the presence of other category (expecting ‘No’). A competent model must be able to identify not only that the object in all images (per category) is the target category but also that it is not some other category. This results in two types of measures. First, for each question type, we measure the average percentage of cases the model correctly answers the question correctly for all images for a given category. We report this accuracy measure per question-type. Additionally, we also measure the average proportion of time the model correctly answers all four questions for all five images in the given category—i.e., all 20 questions associated with a given category. We refer to this latter measure as ‘Joint’ accuracy. Table˜1 shows the accuracy measure across all question types as well as the Joint accuracy.
@@ 4.1 Data
Due to a lack of existing datasets that test knowledge of quantifiers in VLMs in a controlled manner, we design our own Vision and Language dataset, inspired by Gelman et al. (2002) and Smith (1980). To this end, we prompt Gemini 2.5 Flash Image (Comanici et al., 2025) to first generate an image with a specified property that holds true for either all or some objects (e.g., flowers in a vase), we then prompt it to edit the image by adding new objects such that the property now applied to the opposite quantifier (e.g., adding a few flowers outside the vase). This gives us our input context variation between all and some. To ensure the generated images are consistent with our instructions, we manually analyzed every generation. Specifically, we prompted the model to generate 200 pairs of images, and then checked if 1) the pairs were minimally different (no visible difference in backgrounds), still maintained the all vs. some distinction, and 3) did not add additional properties beyond what was specified. This resulted in 140 pairs in total. Appendix C shows full details about our manual annotation. We then paired these images with their respective questions, as described in the previous paragraph (6 different conditions), giving us a total of 840 stimuli. These stimuli mimic those used by Gelman et al. (2002).
@@
Our language stimuli include synthetically generated contexts posed as word-problems about objects and their properties. We generate 150 unique contexts differing in objects (N=5, e.g., coins in a jar, bottles in a crate, etc.), property pairs (N=6, e.g., red/blue in color, contains milk/soda, etc.), and amounts (N=5, e.g., 500 total objects, with 300 having a given property and 200 having the other property). This results in a total of 900 stimuli (when combined with 6 different conditions). The example in Table 2 uses as its object ‘blocks on a table’, as its property pair, ‘red/blue in color’ and as its amount, 20 total objects—12 having the majority property (red) and 8 having the minority property (blue).
@@ C.1 Prompts
In the third condition, we first prompt the model to generate an image with most of the objects (majority) in a container, and some of them (minority) out of it. Then we prompt the model to modify the image by removing the objects that are out of the container.
@@ C.2 Annotation
We manually annotate the 200 generated image pairs using the frontend shown in Figure 6. We excluded 60 pairs from our dataset. Examples of excluded pairs are shown in Figure 7.